OpenStudy (anonymous):

To increase power, it is necessary to decrease the work being done and decrease the time in which the work was completed. increase the work being done or decrease the time in which the work is completed. increase the work being done and increase the time in which the work was completed. decrease the work being done and increase the time in which the work is completed.

OpenStudy (theeric):

Do you know what power is? I'd rather work through this with you than simply answer it. I hope you don't mind.

OpenStudy (anonymous):

sure.

OpenStudy (theeric):

Alright! So, do you know what power is? In physics?

OpenStudy (anonymous):

work being done.

OpenStudy (theeric):

Sort of, and very close. Just missing one thing. Power is also associated with \(time\). Work done in a certain time.

OpenStudy (theeric):

And, if visualizing equations helps you, here you go!\[\large P=\frac{W}{t}\]\[\Huge P=\frac{W}{\large t}\]\[\small{P}=\frac{\large W}{\Huge t}\]

OpenStudy (theeric):

But in general, more work makes power greater, and more time in which the work is done makes power lesser.

OpenStudy (theeric):

This points to one answer.

OpenStudy (theeric):

Any thoughts?

OpenStudy (anonymous):

So it would be the 2nd one?

OpenStudy (theeric):

Yup! :) Congrats!